OpenStudy (anonymous):

What is the solution of the equation shown? -4w+5/3=-13

OpenStudy (whpalmer4):

Is that \[\frac{-4w+5}{3} = -13\]or\[-4w+\frac{5}{3} = -13\]? Please use the equation editor or parentheses so we don't have to ask.

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

Sorry

OpenStudy (whpalmer4):

Remember, the order of operations would suggest the second one, that's why you need the parentheses! Okay, what happens after you multiply both sides of the equation by 3?

OpenStudy (whpalmer4):

You can do just about anything, so long as you do it to both sides.

OpenStudy (anonymous):

The equation would then be -4w+5=-39. Then you subtract 5 from each side, right

OpenStudy (whpalmer4):

That would be good.

OpenStudy (whpalmer4):

I personally would add 4w to both sides and add 39 to both sides, giving me 39+5 = 4w

OpenStudy (anonymous):

From there it would be -4w=-44 which w= 11

OpenStudy (whpalmer4):

I find I'm more likely to make a mistake manipulating lots of negative numbers, so if it is easy to make everything positive, I do so.

OpenStudy (whpalmer4):

and we both get the same answer, of course. Checking the work is important: \[\frac{-4(11)+5}{3} = -13\]\[-44+5 = -13*3\]\[-39=-39\checkmark\]

OpenStudy (anonymous):

Thanks being a homeschool mom is hard with a smart kid.

OpenStudy (whpalmer4):

Probably isn't a picnic for the kid either :-)

OpenStudy (anonymous):

Oh she's fine with it.

OpenStudy (whpalmer4):

How old is your daughter?

OpenStudy (whpalmer4):

I'm not home-schooling my 9-year-old son, but we spend a lot of time doing math together.

OpenStudy (anonymous):

11

OpenStudy (anonymous):

She is in The Gifted program

OpenStudy (whpalmer4):

Here's a little puzzle for her (my son, much to my surprise, did it in his head in a few minutes!): consider all of the possible 3 digit numbers. List all of the numbers such that the sum of the digits * 11 gives you the number.

OpenStudy (anonymous):

She got it right I am just the one who has to check it

OpenStudy (whpalmer4):

If you get stumped on something, feel free to send me a message.

OpenStudy (anonymous):

Do you want to exchange emails